11x+4=6x+29 wats x
Answer by checkley71(8403)

11X+4=6X+29 OR 11X-6X=29-4 OR 5X=25 OR X=25/5 OR X=5
PROOF 11*5+4=6*5+29 OR 55+4=30+29 OR 59=59